OpenStudy (anonymous):

If a die is rolled 1 time then the probability of getting 1)2 or 5 is ? 2) that, 6 does not occur is?

OpenStudy (unklerhaukus):

how many faces does a die (cube ) have?

OpenStudy (anonymous):

Sample Space = 6

OpenStudy (unklerhaukus):

right

OpenStudy (unklerhaukus):

so what is the probability of a rolling a 2?

OpenStudy (anonymous):

1/3

OpenStudy (unklerhaukus):

how many 2's are there on a die?

OpenStudy (anonymous):

i wrote the probability of getting 2 or 5

OpenStudy (anonymous):

1/6

OpenStudy (anonymous):

what to do in 2nd part

OpenStudy (unklerhaukus):

and the probability of a 5 is also 1/6 so adding these you got 1/3, right

OpenStudy (anonymous):

a)1/6,1/6 b)5/6 not confirm.

OpenStudy (unklerhaukus):

well for the second part, se the fact that the sum of probabilities is one, and simply take away the probability of 6, (also 1/6),,, so P(¬6)=1-1/6

OpenStudy (anonymous):

what ? i did'nt understand

OpenStudy (unklerhaukus):

which bit?

OpenStudy (anonymous):

what will be event in which 6 does not occur ?

OpenStudy (anonymous):

u there ?

OpenStudy (unklerhaukus):

well the probability 6 does occur is 1/6, right? and the total probability is 1,

OpenStudy (anonymous):

how is it 1?

OpenStudy (unklerhaukus):

1/6+1/6+1/6+1/6+1/6+1/6=6/6=1

OpenStudy (anonymous):

so is 1/6 in which 6 does not occur?

OpenStudy (unklerhaukus):

1-1/6

OpenStudy (anonymous):

thanks :)

OpenStudy (unklerhaukus):

does it make sense now?